The Taxpayer Assistance Center (TAC) provides taxpayer assistance to help in the preparation of individual tax returns. The amount of assistance (A) that can be provided is a function of the number of professional accountants (P) and trained tax preparers (T) that are employed. The function is given by: A = 500P +200T +50PT

(a) TAC has budgeted $700,000 for a new office to pay for the accountants and the tax preparers. It pays an accountant $100,000 and a tax preparer $50,000. Using the Lagrangean multiplier approach calculate the optimal staff combination that can maximize the amount of tax assistance. Determine the optimal amount of assistance provided by TAC. Explain your calculations.

(b) Using Excel-Solver verify your answer to (a). (Note: Provide the screenshots of the spreadsheet along with Solver windows as shown in the Solver supplement so that all Solver commands can be seen.)
Answer #1

(a)

The amount of assitence is given as

A = 500P 2007, + 507

The budget constraint of the firm is given as

100P50T700

2P+T=14

The firms objective is to

maximize A = 500P 2007, + 507

subject to 2P+T=14

The required Lagrange function is given as

z=left (500P+200T+50PT ight )+lambda (14-2P-T)

The FOCs are

500+ ...................(1)

Öz T200...............(2)

= 14-2P-T......................................(3)

From (1) and (2)

10+T

herefore 2P-T=2.................(4)

Solving (3) and (4):

P=4 ; T=6

================================================================================

(b)

At equilibrium

T=2P-2

herefore A=500P+200(2P-2)+50Pleft ( 2P-2 ight )

A 500P 400P-400 100P2-100P-100P2 + 800P-400

herefore rac{A}{100}=P^2+8P-4=P^2+8P+16-20=(P+4)^2-20

00T

herefore rac{sqrt{A-2000}}{10}=P+4

VA - 2000 10

herefore T=rac{sqrt{A-2000}}{5}-10

C-100000P+50000T-100000 VA-2000 VA - 2000

herefore MC=50000left (rac{1}{10sqrt{A-2000}} ight )+25000left (rac{1}{5sqrt{A-2000}} ight )

( ) +5000( )= 0000 . MC 5000 VA-2000 VA-2000 VA -

This implies the firm has increasing marginal cost.

====================================================================================

(c)

A1 f Microsoft Excel 14.0 Answer Report 1 Microsoft Excel 14.0 Answer Report 2 Worksheet: [Book1]Sheet1 3 Report Created: 5/1

B3 f100000 B12+50000*B13 Solver Parameters 3 outputk...i- 4400 Set Objective: 5 Decision O Max O Min Value Of By Changing Var